OpenStudy (anonymous):

2 square root 3 * 5 square root 27

OpenStudy (anonymous):

help me plz :)

OpenStudy (anonymous):

\[2\sqrt{3}*5\sqrt{27}\]

OpenStudy (anonymous):

= \[2 x \sqrt{3} x 5 x 3\sqrt{3}\]

OpenStudy (anonymous):

because \[\sqrt{27} = \sqrt{3 x 3 x 3} = 3\sqrt{3}\]

OpenStudy (anonymous):

= \[2 x 5 x 3 x (\sqrt{3})^{2} = 2 x 5 x 3 x 3 = 90\]

OpenStudy (anonymous):

90